Fairchild Company has expected earnings of $986,000 and a market value of equity of $11,500,000. The firm is planning to issue $4,880,000 of debt at 6.2 percent interest and use the proceeds to repurchase shares at their current market value. Ignore taxes. What will be the cost of equity after the repurchase?

10.32%

10.87%

11.29%

11.65%

11.91%
Answer #1

Using Modigliani-Miller theorem (without tax effect)

Levered cost of equity = unleveled cost + debt equity ratio (unlevered cost of equity - cost of debt)

Unlevered cost of equity = earnings / value of equity =986,000/ 11,500,000= 8.57%

No tax, so no effect on value of levered firm.

Debt equity ratio of levered firm = debt issue / (value of equity - debt issue)

Debt equity ratio of levered firm = 4,880,000 /(11,500,000-4,880,000)= 0.737

Cost of debt = 6.2%

Levered cost of equity= 8.57% + 0.737 (8.57%- 6.2%)

Levered cost of equity= 8.57% + 1.74669%=10.32%